OpenStudy (anonymous):

two dice are rolled, what is p(sum of 7) s=36

OpenStudy (raden):

the events have sum 7 are (1,6),(6,1),(2,5),(5,2),(4,3),(3,4) so, there are 6 events. P(sum of 7) = 6/36 = 1/6

OpenStudy (anonymous):

this is the same answer i got, and its telling me that its wrong so i dont understand maybe i should round up to 0.17, instead of entering 0.16